OpenStudy (hitaro9):

From the problem I obtained the equation "1800(2^(x/40))." I need to convert that to an equation that looks like this A(e^kx). What do A and k need to be for that to be true? No idea how to do this. Can someone help?

OpenStudy (anonymous):

\[\log_b x^n = n \log_b x\]I think that should help you.

OpenStudy (hitaro9):

I'm honestly still lost. What happens when I take the ln of 1800(2^x/40)?

OpenStudy (hitaro9):

Does it become x/40 (ln(1800(2)) ?

OpenStudy (anonymous):

I misread it. It's not that complicated. Sorry.

OpenStudy (hitaro9):

Okay lol. I still don't completely understand the whole e thing and how it relates to exponential growth

OpenStudy (anonymous):

Essentially what you have is \(2^{\frac{1}{40}x} = e^{kx}\) You can just say A = 1800 and stick that back in later.

OpenStudy (hitaro9):

Okay. So how do I get k then?

OpenStudy (hitaro9):

Thanks so much for the help btw.

OpenStudy (anonymous):

I think you just have to think of it as \[e^{ln 2^{\frac{1}{40}}x}\]

OpenStudy (anonymous):

That 1/40 is the exponent of the 2, by the way. It's hard to tell in that tiny font.

OpenStudy (hitaro9):

Okay. I see. So my final answer would be 1800(e^xln2^1/40)

OpenStudy (anonymous):

I am not 100% sure about this! But I think that's how you convert bases like that. Like any number can be written as \(n = e^{\ln n}\)

OpenStudy (hitaro9):

Okay. I do have 3 guesses on my online assignment, so if it ends up not being right I'll let you know. 1 sec

OpenStudy (hitaro9):

So A was right, it is 1800, but k was not. So yeah, we are left with how to convert 2^(x/40) to e^kx

OpenStudy (anonymous):

Ok, hold on. I think I get it now.

OpenStudy (anonymous):

\[2^{\frac{1}{40}x} = e^{kx}\]\[\frac{1}{40}x = \log_2 e^{kx}\]\[\frac{1}{40}x = \frac{\ln e^{kx}}{\ln 2}\]\[\frac{1}{40}x =\frac{kx}{\ln 2}\]\[\frac{\ln 2}{40}x = kx\]\[\frac{\ln 2}{40} = k\]

OpenStudy (anonymous):

Tell me if that makes sense. :D

OpenStudy (hitaro9):

That makes a lot of sense actually.

OpenStudy (anonymous):

I think the first time I tried to shortcut doing all that math and bungled something somewhere. But that one should work. And I can actually even explain it now, if needed. :P

OpenStudy (hitaro9):

It was the wrong answer, even though I'm pretty sure it's right :/ Ugh. Okay, it's fine, I'm pretty sure I understand how you did that. I appreciate the help.

OpenStudy (anonymous):

Ok, well, I don't know what they're expecting, then. :( Sorry.

OpenStudy (hitaro9):

I have 1 guess left, but I'm not sure how I should enter it. Like, I'm fairly certain the answer you gave is right, cause I actually understand how you got it and it makes sense to me

OpenStudy (anonymous):

Also, I just noticed, I had it right before, too, just written in a weird form. But they are equal because of the very first thing I said. :D

OpenStudy (hitaro9):

Ah yeah I see.

OpenStudy (anonymous):

\[\frac{1}{40} \times \ln 2 = \ln 2^{\frac{1}{40}}\] Now I'm double convinced!

OpenStudy (hitaro9):

Alright. I'll try to figure out how they want it entered, I really appreciate the help dude

OpenStudy (hitaro9):

I understand it a lot better now, thanks again.